OpenStudy (anonymous):

There are three terms in Arithmetic Sequence . The common difference is 11 . A1=n-5 , A2=n-1 , and A3=2n . What are the terms ?

hartnn (hartnn):

common difference = A2-A1 = A3-A2 = 11

hartnn (hartnn):

A3-A2 = 11 2n - (n-1) =11 can you find 'n' from here ??

OpenStudy (anonymous):

Hmm ...

hartnn (hartnn):

was that a yes or no ?? O.o

OpenStudy (anonymous):

Wait .

hartnn (hartnn):

waiting...

OpenStudy (agent0smith):

A2 - A1 = 11 (n - 1) - (n -5) = 11 -1 + 5 = 11 no. 4 does not equal 11.

hartnn (hartnn):

ok...i guess, 11 as common difference(d) should naot have been given. in that case d = A2 - A1 = 4 then A3-A2 = 4 and find 'n'.

OpenStudy (anonymous):

How can I find the n ?

OpenStudy (anonymous):

why 4 again??

hartnn (hartnn):

common diff. always remains constant/same!

hartnn (hartnn):

2n - (n-1) =4

hartnn (hartnn):

n+1=4

OpenStudy (anonymous):

n=5??

hartnn (hartnn):

is 5+1 = 4 ??

OpenStudy (anonymous):

n=3 ?

hartnn (hartnn):

yes. n=3 :)

OpenStudy (agent0smith):

Yep, common difference of 4 works here.

OpenStudy (anonymous):

n-5=-2 n-1=2 2n=6

hartnn (hartnn):

\(\checkmark \)

OpenStudy (anonymous):

haha i used this d = A2 - A1 = 4

OpenStudy (agent0smith):

Note that we can't guarantee these are the correct answers, since... the question claims the difference is 11. Either that's wrong, or you didn't write out the question correctly.

OpenStudy (anonymous):

That's okay :))

OpenStudy (anonymous):

WHAT IF THE COMMON DIFFERENCE IS 9??

hartnn (hartnn):

from d = A2 - A1 the common diff. must be 4 cannot be anything else.

OpenStudy (anonymous):

OH WELL. xD

hartnn (hartnn):

did you ppl start learning sequences ??

OpenStudy (anonymous):

yup,xD arithmetic seq and series and mean and geom seq and series

OpenStudy (anonymous):

the formulae actually are mixing in my head. O_o

hartnn (hartnn):

then going through this might be helpful.... http://openstudy.com/users/hartnn#/updates/503bb2a0e4b007f9003103b0

OpenStudy (agent0smith):

\[\large{\color{Blue}{T}\color{Yellow}{U}\color{Magenta}{T}\color{Pink}{O}\color{MidnightBlue}{R} I\color{Blue}{A}\color{Green}{L}!}\] ^^ Is it a tutorial on how to have a seizure? :P

OpenStudy (amistre64):

nice "tutorial" :)